OpenStudy (anonymous):

simplify sqrt(1-a)+h *(sqrt(1-a))

OpenStudy (anonymous):

Is this the equation?: \[\sqrt{1-a}+h\sqrt{1-a}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{1-a}(1+h)\]

OpenStudy (anonymous):

Alright. So the principle here is factoring. Watch carefully: Let \(u=\sqrt{1-a}\). We can simplify this to \[\sqrt{1-a}+h\sqrt{1-a}=u+hu.\] YOO-HOO! lol. Sorry. minor observation. ehh so normally if we have a formula in the form: \(a(1+b)=a+ab\). In this case \(u+hu=u(1+h)\). So then we can substitute \(\sqrt{1-a}\) back in: \[u+hu=u(1+h)=\sqrt{1-a}(1+h)\] as authentication error aptly "blurted" haha ;)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Anytime!